# Emergent Hawking Radiation and Quantum Sensing in a Quenched Chiral Spin Chain

### v1

@@ B.2 Gaussian Wave Packets
This ratio is, in general, non-thermal. This deviation from perfect thermality can be attributed to grey-body factors Unruh [1976a], Page [1976], which encode the fact that modes generated near the horizon do not propagate to the detector without modification. As they travel through the background geometry, portions of the wave packet are partially reflected, scattered, or otherwise distorted, leading to a frequency-dependent filtering of the detected spectrum. Furthermore, for low-frequency outgoing modes ($f\simeq 0$), the ratio reduces to $\frac{\alpha_{ff^{\prime}}^{G}}{\beta_{ff^{\prime}}^{G}}\approx 1$, indicating an equal contribution from positive and negative-frequency components. As a consequence, the particle occupation number is given by
@@ C.1 Exact Dynamics in Strong Coupling
where $P_{ee}(p,t)=1-P_{gg}(p,t)$ and $P_{ge}(p,t)=P_{eg}^{\ast}(p,t)$. The Rabi frequency is defined as $\Omega_{R}=2\sqrt{g_{1}^{2}+g_{2}^{2}}\sin{2p}$. The total population is the product over all quasi-momenta $p$. The decoherence function Liu et al. [2009], defined as $D(t)=\prod_{p}|P_{ge}(p,t)|^{2}$, quantifies the magnitude of coherence present in the reduced density matrix. We emphasize that since the qubit starts in an energy eigenstate ($D(0)=0$ by construction), $D(t)$ measures the suppression of dynamically generated coherence. As the qubit-chiral interaction is switched on, $D(t)$ initially increases, indicating the buildup of coherence. At later times, however, it saturates, reflecting the qubit completely entangled with the environmental degrees of freedom. This definition differs from commonly used decoherence measures, where $D(t)=1$ corresponds to a pure, fully coherent state and $D(t)=0$ signals complete loss of coherence.
@@ C.2 Weak Coupling Rates
where $f_{D}(\omega)=\left(e^{\omega/T}+1\right)^{-1}$. These rates yield the population dynamics:
@@
At the horizon ($|\mathcal{U}|=v$), the bath spectral density diverges ($J(\omega)\rightarrow\infty$), signaling infinitely strong mode coupling. In this limit, the qubit states are fully governed by the Dirac thermal spectrum. Thus, as long as the analog black hole exists, the detector couples to the Dirac spectrum rather than the vacuum, allowing $T_{H}$ to be extracted as an emergent property of the environment.

### v2

@@ B.2 Gaussian Wave Packets
This ratio is, in general, non-thermal. This deviation from perfect thermality can be attributed to grey-body factors Unruh (1976); Page (1976), which encode the fact that modes generated near the horizon do not propagate to the detector without modification. As they travel through the background geometry, portions of the wave packet are partially reflected, scattered, or otherwise distorted, leading to a frequency-dependent filtering of the detected spectrum. Furthermore, for low-frequency outgoing modes ($f\simeq 0$), the ratio reduces to $\frac{\alpha_{ff^{\prime}}^{G}}{\beta_{ff^{\prime}}^{G}}\approx 1$, indicating an equal contribution from positive and negative-frequency components. As a consequence, the particle occupation number is given by
@@ C.1 Exact Dynamics in Strong Coupling
where $P_{ee}(p,t)=1-P_{gg}(p,t)$ and $P_{ge}(p,t)=P_{eg}^{\ast}(p,t)$. The Rabi frequency is defined as $\Omega_{R}=2\sqrt{g_{1}^{2}+g_{2}^{2}}\sin{2p}$. The total population is the product over all quasi-momenta $p$. The decoherence function Liu et al. (2009), defined as $D(t)=\prod_{p}|P_{ge}(p,t)|^{2}$, quantifies the magnitude of coherence present in the reduced density matrix. We emphasize that since the qubit starts in an energy eigenstate ($D(0)=0$ by construction), $D(t)$ measures the suppression of dynamically generated coherence. As the qubit–chiral interaction is switched on, $D(t)$ initially increases, indicating the buildup of coherence. At later times, however, it saturates, reflecting the qubit completely entangled with the environmental degrees of freedom. This definition differs from commonly used decoherence measures, where $D(t)=1$ corresponds to a pure, fully coherent state and $D(t)=0$ signals complete loss of coherence.
@@ C.2 Weak Coupling Rates
where $f_{D}(\omega)=\left(e^{\omega/T}+1\right)^{-1}$. These rates yield the population dynamics of the excited state:
@@
At the horizon ($|\mathcal{U}|=|v|$), the bath spectral density diverges ($J(\omega)\rightarrow\infty$), signaling infinitely strong mode coupling. In this limit, the qubit states are fully governed by the Dirac thermal spectrum. Thus, as long as the analog black hole exists, the detector couples to the Dirac spectrum rather than the vacuum, allowing $T_{H}$ to be extracted as an emergent property of the environment.